OpenStudy (theharleysmith):

Marge cut 16 pieces of tape. Each piece is 3/8in long HOW MUCH TAPE DID SHE USE

OpenStudy (theharleysmith):

@codyhelton75 Help please

OpenStudy (kendricklamar2014):

This is pretty simple to solve actaully. All you need to do is: \[16 \times \frac{ 3 }{ 8 }\] \[= \frac{ 16 }{ 1 } \times \frac{ 3 }{ 8 }\] \[= \frac{ 48 }{ 8 }\] \[\frac{ 48 }{ 8 } = 6\] Answer: 6